# Optimization of a cosmic muon tomography scanner for cargo border control inspection

### v1

@@ VI.2 Optimization with BCA inference method
FIG. 21 presents the convergence of the best minimum objective, showing a reduction throughout the optimization.
@@ VI.3 Optimization with inference method
In this approach, BO was performed using the inference method, where the evaluation function was defined as the mean squared error (MSE) between normalized inferred and ground truth values. Normalization was necessary due to the significant underbias observed for low-Z materials, particularly air (FIG. 19(b)). Since air is present in the passive volume, the raw MSE values would otherwise be orders of magnitude larger than what is expected from a typical MSE evaluation, potentially distorting the optimization process. FIG. 25 presents the convergence of the best minimum objective, showing a reduction throughout the optimization. FIG. 26 illustrates the response surface for the mean and standard error of the surrogate objective. Unlike the case with BCA inference, the gap sizes between the upper and lower hodoscope pairs (FIG. 27) do not exhibit a clear optimization trend, with the best-evaluated parameters appearing mid-range in the trials and the observed mean of the objective function remaining relatively stable throughout the GPEI stage following Sobol sampling. This behavior can be attributed to the noisiness of the MSE, which arises from the inherent bias in the inference method.
@@
To statistically compare the inference performance of the optimized detector configurations, we evaluate the Mean Squared Error (MSE) and the Structural Similarity Index Measure (SSIM) for 100 simulated passive volumes, with and without a high-density uranium block (FIG. 29). While the MSE directly quantifies voxel-wise differences, the SSIM is a perceptual metric that evaluates the structural similarity between predicted and true volumes. It is computed over local regions using a sliding window of size , and the final SSIM score is obtained by averaging over all such windows. The SSIM is defined as:
SSIM | ||||
@@ VII Conclusions
We have presented the current status of ongoing studies toward the optimization of a future SilentBorder scanner for muon scattering tomography. Studies using the GEANT4 framework have shown that minimizing the in-plane gaps between active hodoscope areas, by keeping any casing and electronics sizes as small as possible, will ensure optimal efficiency of the system. However, increasing the vertical spacing between hodoscope plates does show improvement in the angular resolution with little reduction in efficiency and is worth pursuing. A complementary GEANT4 study on the role of secondary particle hits in event reconstruction was also conducted. Our findings suggest that secondary particles do not provide any extra material discrimination power in a ‘SilentBorder’ type scanner but also cause very limited degradation of the reconstruction quality and therefore do not need separate measures to exclude them.
@@ Acknowledgements.
The authors acknowledge funding from the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”) and by the Fonds de la Recherche Scientifique - FNRS under Grants No. T.0099.19 and J.0070.21. Pietro Vischia’s work was supported by the “Ramón y Cajal” program under Project No. RYC2021-033305-I funded by the MCIN MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR.Data Availability

### v2

@@ VI.2 Optimization with BCA inference method
Fig. 20 shows the convergence of the best minimum objective over 100 trials on CPU, showing a 13.3% reduction throughout the optimization. For simplicity, only a single run is shown, although different random seeds may yield different results. The full optimization takes 26 minutes, while using gradient descent would result in 38% increase in runtime. This difference arises because the BCA reconstruction is computationally heavier than the formula-based inference method, involving PoCA pair-wise voxel clustering. The gradient descent optimization requires evaluating many gradients per step, resulting in longer runtime. With BO, the GP model reduces the number of expensive evaluations, so it is faster overall despite the overhead of building the GP model.
@@ VI.3 Optimization with inference method
In this approach, BO was performed using the inference method, where the evaluation function was defined as the mean squared error (MSE) between normalized inferred and ground truth values. Normalization was necessary due to the significant underbias observed for low-Z materials, particularly air (Fig. 18(b)). Since air is present in the passive volume, the raw MSE values would otherwise be orders of magnitude larger than what is expected from a typical MSE evaluation, potentially distorting the optimization process. Fig. 24 shows the convergence of the best minimum objective, with a total reduction of 22.8%. The complete optimization runs 11 minutes on CPU with BO, comparable to 9 minutes with gradient descent. Here, the formula-based inference method is relatively computationally inexpensive, so the additional runtime in BO stems primarily from the overhead of updating the GP surrogate model.
Fig. 25 illustrates the response surface for the mean and standard error of the surrogate objective. Unlike the case with BCA inference, the gap sizes between the upper and lower hodoscope pairs (Fig. 26) do not exhibit a clear optimization trend, with the best-evaluated parameters appearing mid-range in the trials and the observed mean of the objective function remaining relatively stable throughout the GPEI stage following Sobol sampling. This behavior can be attributed to the noisiness of the MSE, which arises from the inherent bias in the inference method.
@@
To statistically compare the inference performance of the optimized detector configurations, we evaluate the Mean Squared Error (MSE) and the Structural Similarity Index Measure (SSIM) for 100 simulated passive volumes, with and without a high-density uranium block (Fig. 28). The performance is compared with a baseline corresponding to a detector design based on human intuition for maximizing acceptance (Fig. 29). While the MSE directly quantifies voxel-wise differences, the SSIM is a perceptual metric that evaluates the structural similarity between predicted and true volumes. It is computed over local regions using a sliding window of size , and the final SSIM score is obtained by averaging over all such windows. The SSIM is defined as:
SSIM(infer, true) =
(2μinferμtrue+ C1)(2σinfer, true+ C2)(μinfer2+ μtrue2+ C1)(σinfer2+ σtrue2+ C2)
@@
Some of the remaining limitations of TomOpt arise from intrinsic challenges in muography: the inverse problem is inherently ill-posed and relies on approximations, with the PoCA method reducing multiple scattering to a single point and introducing systematic biases. Importantly, nearly all current reconstruction methods build on PoCA, meaning that the sensitivity of detector design to volume reconstruction is closely tied to the accuracy of this approximation, highlighting the importance of developing approaches that go beyond PoCA in future work.
@@ VII Conclusions
This work outlines a clear pathway toward the optimization of a future SilentBorder scanner for muon scattering tomography, with current results demonstrating meaningful advances in the studied approaches. Future studies will aim to extend these developments toward more realistic operational scenarios and validation with experimental data. Studies using the GEANT4 framework have shown that minimizing the in-plane gaps between active hodoscope areas, by keeping any casing and electronics sizes as small as possible, will ensure optimal efficiency of the system. However, increasing the vertical spacing between hodoscope plates does show improvement in the angular resolution with little reduction in efficiency and is worth pursuing. A complementary GEANT4 study on the role of secondary particle hits in event reconstruction was also conducted. Our findings suggest that secondary particles do not provide any extra material discrimination power in a ‘SilentBorder’ type scanner but also cause very limited degradation of the reconstruction quality and therefore do not need separate measures to exclude them.
@@ Acknowledgements.
The authors acknowledge funding from the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”) and by the Fonds de la Recherche Scientifique - FNRS under Grants No. T.0099.19 and J.0070.21. Pietro Vischia’s work was supported by the “Ramón y Cajal” program under Project No. RYC2021-033305-I funded by the MCIN MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. Finally, the authors gratefully acknowledge the computer resources at Artemisa, funded by the European Union ERDF and Comunitat Valenciana as well as the technical support provided by the Instituto de Fisica Corpuscular, IFIC (CSICUV).Data Availability